The Sum of the digits of a two-digit number is 10. The ones digit is six more than the tens digit. What is the number ?
Answer by oberobic(2304)

With sum of digits problems, you have to keep in mind that the variables we use are "standing" beside each other, not multiplying. So, if we define the two-digit number as 'xy' then we mean 'x' is standing next to 'y', not that we are multiplying x*y.
.
xy has the value 10*x + y because the 'x' is in the tens place and 'y' is in the ones place.
.
We are told:
.
x+y = 10
y = x+6
.
Substitute y = x+6
.
x + x+6 = 10
2x = 4
x = 2
.
y = x+6
y = 8
.
The number is 28.
.
Done.